# A First-order Generative Bilevel Optimization Framework for Diffusion Models

### v1

@@ Abstract
Diffusion models, which iteratively denoise data samples to synthesize high-quality outputs, have achieved empirical success across domains. However, optimizing these models for downstream tasks often involves nested bilevel structures—such as tuning hyperparameters for fine-tuning tasks or noise schedules in training dynamics—where traditional bilevel methods fail due to the infinite-dimensional probability space and prohibitive sampling costs. We formalize this challenge as a generative bilevel optimization problem and address two key scenarios: (1) fine-tuning pre-trained models via an inference-only lower-level solver paired with a sample-efficient gradient estimator for the upper level, and (2) training diffusion model from scratch with noise schedule optimization by reparameterizing the lower-level problem and designing a computationally tractable gradient estimator. Our first-order bilevel framework overcomes the incompatibility of conventional bilevel methods with diffusion processes, offering theoretical grounding and computational practicality. Experiments demonstrate that our method outperforms existing fine-tuning and hyperparameter search baselines.
@@ 1 Introduction
Bilevel optimization, which involves nested problems where a lower-level optimization is constrained by the solution of an upper-level objective, has evolved from its theoretical origins in the 1970s (Bracken & McGill, 1973) into a cornerstone of modern machine learning. Its ability to model hierarchical dependencies makes it ideal for complex learning tasks, such as hyperparameter tuning (Pedregosa, 2016), meta-learning (Finn et al., 2017), reinforcement learning (Stadie et al., 2020; Shen et al., 2024b), adversarial training (Zhang et al., 2022), neural architecture search (Liu et al., 2019) and LLM alignment (Zakarias et al., 2024; Shen et al., 2024a; Gong et al., 2022; Qin et al., 2024).
@@
The key challenges of solving the generative bilevel problem (1) are threefold. First, the lower-level variable $y$ is typically a distribution that operates in infinite dimensional probability space. However, direct access to or optimization over distribution is not feasible; we only have access to samples, and estimating distributions from samples is highly sample-inefficient. Therefore, the gradient over the distribution is inaccessible, making traditional gradient-based bilevel approaches (Shen et al., 2023; Kwon et al., 2023; Ji et al., 2021; Chen et al., 2021; Hong et al., 2023; Ghadimi & Wang, 2018) generally inapplicable. Second, the objectives $f(x,y)$ and $g(x,y)$ are usually some measures of the sample quality and might not have an explicit form in terms of the hyperparameter $x$, so that $\nabla_{x}f(x,y)$ and $\nabla_{x}g(x,y)$ are either not explicitly given or requires sample efficient approximation. Third, for fine-tuning the diffusion model, existing literature related to bilevel fine-tuning on diffusion model (Clark et al., 2024; Marion et al., 2024) requires backpropagations over the pre-trained model, which often suffers from high computational and memory costs. In contrast, we design an inference-only bilevel method for this task.
@@ 2.1 Bilevel optimization
An efficient approach to solving bilevel optimization in (1) is to reformulate (1) to its single-level penalty problem (Shen et al., 2023; Kwon et al., 2024), given by
@@
and then optimize the upper-level and lower-level variables jointly. By setting the penalty constant $\gamma={\cal O}(\epsilon^{-0.5})$ inversely proportional to the target accuracy $\epsilon$, the single-level problem (2) is an ${\cal O}(\epsilon)$ approximate problem to the original bilevel problem (1). This method builds upon equilibrium backpropagation (Scellier & Bengio, 2017; Zucchet & Sacramento, 2022), which introduced a similar framework for strongly convex lower-level problems. However, recent works extend the study to accommodate some nonconvex lower-level problems and lay the foundation for broader applications (Shen et al., 2023; Kwon et al., 2024).
@@
[…]
@@ 2.2 Diffusion models
$\mathrm{d}\widetilde{U}_{t}$ $=\left[\frac{1}{2}q(T-t)\widetilde{U}_{t}+{\nabla\log p_{T-t}\left(\widetilde{U}_{t}\right)}\right]\mathrm{d}t$
$+\sqrt{q(T-t)}\mathrm{d}\bar{W}_{t},~{}~{}~{}\text{ for }~{}t\in[0,T),$ (6)
where $p_{t}(\cdot)$ is the marginal distribution of $U_{t}$ in the forward process. Let $u$ denote the realization of a random variable $U_{t}$ and $s(u,t):=\nabla\log p_{t}(u)$ is the unknown score function that has to be estimated in practice.
@@
$\mathrm{d}\widetilde{U}_{t}$ $=\left[\frac{1}{2}q(T-t)\widetilde{U}_{t}+s_{\theta}(\widetilde{U}_{t},T-t)+\textit{G}(\widetilde{U}_{t},T-t)\right]\mathrm{d}t$
$+\sqrt{q(T-t)}\mathrm{d}\bar{W}_{t},~{}~{}t\in[0,T).$ (8)
@@ 3.1 Reward fine-tuning in diffusion models
[…]
@@ 3.2 Noise scheduling in diffusion models
[…]
@@ 4.1 Strategy with pre-trained diffusion models
where $p_{\gamma}^{*}(\lambda)\in\operatorname*{arg\,min}_{p}\mathcal{L}_{\gamma}(\lambda,p)$ and $p^{*}(\lambda)\in\operatorname*{arg\,min}_{p}g(\lambda,p)$. Following the gradient-based bilevel method (Kwon et al., 2023; Shen et al., 2023), a direct way to estimate $\nabla\mathcal{L}_{\gamma}^{*}(\lambda)$ is to use guided backward sampling in Algorithm 5 with (11a) and (11b) to obtain samples from $p^{*}(\lambda)$ and $p_{\gamma}^{*}(\lambda)$, and then compute the KL divergence from samples using kernel-based probability estimation. However, it has two drawbacks: 1) each $\lambda$ update requires guided backward sampling, which is computationally expensive; and 2) when the number of samples is less than the dimensionality of the data, the covariance matrix becomes rank-deficient, which makes kernel-based density estimation impossible.
@@
5:     estimate ¯$\bar{\nabla}\mathcal{L}_{\gamma}^{*}(x_{k})$ with $y^{k}\approx y^{*}$ and $z^{k}\approx z^{*}$ in (4)
@@
4:     estimate ¯$\bar{\nabla}\mathcal{L}_{\gamma_{k}}^{*}(\lambda_{k})$ by (33)
@@ 4.2 Strategy without pre-trained diffusion models
~$\widetilde{\nabla}_{\theta}\mathrm{L}_{\text{SQ}}(\tilde{u}_{\theta,q}^{m})$ $=\frac{\xi}{2\nu}(\mathrm{L}_{\text{SQ}}(\tilde{u}_{\theta+\nu\xi,q}^{m})-\mathrm{L}_{\text{SQ}}(\tilde{u}_{\theta-\nu\xi,q}^{m}))$ (14)
@@
Parametrization for noise scheduler. To further reduce the memory cost and ensure the nondecreasing of $q(t)$, we parameterize the noise scheduler by the commonly used cosine and sigmoid function (Nichol & Dhariwal, 2021; Chen, 2023) and optimize the parameters within these functions instead of directly optimizing $\{q(t)\}_{t=1}^{T}$. In both parametrizations, we optimize just $4$ scalar parameters, significantly reducing the dimensionality of the optimization problem from $T$ to $4$. Specifically, we use
@@
where $q_{s},q_{e},q_{\epsilon},q_{\tau}$ represent the start, end, offset error, and power effect, respectively. Sigmoid parameterization is defined similarly in (34) and can be found in Appendix F.2. Since $q(t)$ should be nondecreasing, we assign $q(t)=1-l(t)/l(t-1)$. After parameterization, ZO perturbation will be added on $q_{s},q_{e},q_{\epsilon},q_{\tau}$ instead of directly on $q(t)$.
@@ 5 Theoretical Guarantee
This theorem demonstrates that when $G_{\eta_{k},\gamma_{k}}(x_{k})\neq 0$, and setting penalty constant $\gamma_{k}\geq\frac{4\sqrt{3}B}{\|G_{\eta_{k},\gamma_{k}}(x_{k})\|}$, the bilevel algorithm will have strict descent over the hyper-function $F(x)$. Otherwise, since ¯$\|\bar{\nabla}\mathcal{L}_{\gamma}^{*}(x)-\nabla F(x)\|={\cal O}(1/\gamma_{k})$, $G_{\eta_{k},\gamma_{k}}(x_{k})=0$ is approximately the stationary points of $F(x)$. The proof can be found in Appendix D.3.
@@
Grid search $125.77\ {\scriptstyle\pm 1.3}$ $31.72\ {\scriptstyle\pm 2.1}$ $3.34$
Random search $123.70\ {\scriptstyle\pm 4.3}$ $35.70\ {\scriptstyle\pm 1.1}$ $3.37$
Bayesian search $140.35\ {\scriptstyle\pm 2.4}$ $33.29\ {\scriptstyle\pm 2.9}$ $12.7$
Weighted sum $116.52\ {\scriptstyle\pm 3.9}$ $36.50\ {\scriptstyle\pm 1.1}$ $4.31$
@@
Table 1: Best FID and CLIP score given by different baselines and our method with penalty constant $\gamma=10^{3}$ for fine-tuning diffusion model application. Running time is measured in hours.
@@ 6 Numerical Experiments
In this section, we present the experimental results of the proposed bilevel-diffusion algorithms in two applications—reward fine-tuning and noise scheduling for diffusion models—and compare them with baseline hyperparameter optimization methods: grid search, random search, and Bayesian search (Snoek et al., 2012).
@@ 6.1 Reward fine-tuning in diffusion models
6.2 Noise scheduling in diffusion models
We evaluated our bilevel noise scheduling method, detailed in Algorithm 6, paired with DDIM backward sampling for the image generation on the MNIST dataset. We trained a U-Net with $178$ layers and $10^{6}+$ parameters following  [^1] https://github.com/bot66/MNISTDiffusion/tree/main. We considered both cosine and sigmoid parametrization and tuned $4$ parameters $q_{s},q_{e},q_{\tau},q_{\epsilon}$ jointly.
Table 2 presents the best FID, inception score (IS), and time complexity achieved by each method. The bilevel method achieves comparable performance with the hyperparameter optimization baselines in both FID and IS while being 6×6$\times$6 × time faster. Generated images by each method are shown in Figure 6. While the Bayesian approach achieves relatively better FID and IS, it tends to focus on generating simpler numbers, such as $1$ and $7$. In contrast, the images produced by the bilevel method show excellent diversity across numbers $0-9$ while maintaining image fidelity.
@@ 7 Related works
Fine-tuning diffusion models. Fine-tuning diffusion models aims to adapt pre-trained models to boost the reward on downstream tasks. Methods in this domain include directly backpropagating the reward (Clark et al., 2024), RL-based fine-tuning (Fan et al., 2024; Black et al., 2023), direct latent optimization (Tang et al., 2024; Wallace et al., 2023), guidance-based approach (Guo et al., 2024; Chung et al., 2022) and optimal control (Uehara et al., 2024). Although entropy regularization is often incorporated into the reward to prevent over-optimization, no existing work has explored designing an efficient bilevel method to tune its strength.
Noise scheduling in diffusion models. Noise schedule is crucial to balance the computational efficiency with data fidelity during image generation. Early works, such as DDPM (Ho et al., 2020), employed simple linear schedules for noise variance, while Nichol & Dhariwal (2021) introduced cosine schedules to enhance performance. Recent studies (Lin et al., 2024; Chen, 2023) have highlighted limitations in traditional noise schedules and proposed new parameterization to improve the image quality. However, none of them considered using bilevel optimization to automatically learn the noise schedule.
Bilevel hyperparameter optimization. Bilevel optimization has been explored as an efficient hyperparameter optimization framework, including hypernetwork search (Mackay et al., 2019; Liu et al., 2019), hyper-representation (Franceschi et al., 2018), regularization learning (Shaban et al., 2019) and data reweighting (Shaban et al., 2019; Franceschi et al., 2017). Recently, it has been explored in federated learning (Tarzanagh et al., 2022) and LLM fine-tuning (Shen et al., 2024a; Zakarias et al., 2024). None of the existing works have explored hyperparameter optimization in diffusion models, and the methods proposed so far are inapplicable due to the infinite-dimensional probability space and the high computational cost of sampling.

### v2

@@ Abstract
Diffusion models, which iteratively denoise data samples to synthesize high-quality outputs, have achieved empirical success across domains. However, optimizing these models for downstream tasks often involves nested bilevel structures, such as tuning hyperparameters for fine-tuning tasks or noise schedules in training dynamics, where traditional bilevel methods fail due to the infinite-dimensional probability space and prohibitive sampling costs. We formalize this challenge as a generative bilevel optimization problem and address two key scenarios: (1) fine-tuning pre-trained models via an inference-only lower-level solver paired with a sample-efficient gradient estimator for the upper level, and (2) training diffusion model from scratch with noise schedule optimization by reparameterizing the lower-level problem and designing a computationally tractable gradient estimator. Our first-order bilevel framework overcomes the incompatibility of conventional bilevel methods with diffusion processes, offering theoretical grounding and computational practicality. Experiments demonstrate that our method outperforms existing fine-tuning and hyperparameter search baselines. Our code has been released at https://github.com/afmsaif/bilevel_diffusion.
@@ 1 Introduction
Bilevel optimization, which involves nested problems where a lower-level optimization is constrained by the solution of an upper-level objective, has evolved from its theoretical origins in the 1970s (Bracken & McGill, 1973) into a cornerstone of modern machine learning. Its ability to model hierarchical dependencies makes it ideal for complex learning tasks, such as hyperparameter tuning (Pedregosa, 2016), meta-learning (Finn et al., 2017), reinforcement learning (Stadie et al., 2020; Shen et al., 2024), adversarial training (Zhang et al., 2022), neural architecture search (Liu et al., 2019) and LLM alignment (Zakarias et al., 2024; Shen et al., 2025a; Gong et al., 2022; Qin et al., 2024).
@@
The key challenges of solving the generative bilevel problem (1) are threefold. First, the lower-level variable $y$ is typically a distribution that operates in infinite dimensional probability space. However, direct access to or optimization over distribution is not feasible; we only have access to samples, and estimating distributions from samples is highly sample-inefficient. Therefore, the gradient over the distribution is inaccessible, making traditional gradient-based bilevel approaches (Shen et al., 2025b; Kwon et al., 2023; Ji et al., 2021; Chen et al., 2021; Hong et al., 2023; Ghadimi & Wang, 2018) generally inapplicable. Second, the objectives $f(x,y)$ and $g(x,y)$ are usually some measures of the sample quality and might not have an explicit form in terms of the hyperparameter $x$, so that $\nabla_{x}f(x,y)$ and $\nabla_{x}g(x,y)$ are either not explicitly given or requires sample efficient approximation. Third, for fine-tuning the diffusion model, existing literature related to bilevel fine-tuning on diffusion model (Clark et al., 2024; Marion et al., 2024) requires backpropagations over the pre-trained model, which often suffers from high computational and memory costs. In contrast, we design an inference-only bilevel method for this task.
@@ 2.1 Bilevel optimization
An efficient approach to solving bilevel optimization in (1) is to reformulate (1) to its single-level penalty problem (Shen et al., 2025b; Kwon et al., 2024), given by
@@
and then optimize the upper-level and lower-level variables jointly. By setting the penalty constant $\gamma={\cal O}(\epsilon^{-0.5})$ inversely proportional to the target accuracy $\epsilon$, the single-level problem (2) is an ${\cal O}(\epsilon)$ approximate problem to the original bilevel problem (1). This method builds upon equilibrium backpropagation (Scellier & Bengio, 2017; Zucchet & Sacramento, 2022), which introduced a similar framework for strongly convex lower-level problems. However, recent works extend the study to accommodate some nonconvex lower-level problems and lay the foundation for broader applications (Shen et al., 2025b; Kwon et al., 2024).
@@
$\text{ and }~~\mathcal{S}_{\gamma}(x):=\operatorname*{arg\,min}_{y\in\mathcal{P}}\mathcal{L}_{\gamma}(x,y).$ (3)
@@ 2.2 Diffusion models
$\mathrm{d}\widetilde{U}_{t}$ $=\left[-\frac{1}{2}q(t)\widetilde{U}_{t}-q(t){\nabla\log p_{t}\left(\widetilde{U}_{t}\right)}\right]\mathrm{d}t$
$+\sqrt{q(t)}\mathrm{d}\widetilde{W}_{t},~~~\text{ for }~t\in(0,T],$ (6)
where $\mathrm{d}\widetilde{W}_{t}$ is the reverse-time Wiener process, $p_{t}(\cdot)$ is the marginal distribution of $U_{t}$ in the forward process. Let $u_{t}$ denote the realization of a random variable $U_{t}$ and $s(u,t):=\nabla\log p_{t}(u)$ is the score function that has to be estimated in practice or given by the pre-trained model.
@@
$\mathrm{d}\widetilde{U}_{t}$ $=\left[-\frac{1}{2}q(t)\widetilde{U}_{t}-q(t)s_{\theta}(\widetilde{U}_{t},t)+\textit{G}(\widetilde{U}_{t},t)\right]\mathrm{d}t$
$+\sqrt{q(t)}\mathrm{d}\widetilde{W}_{t},~~t\in(0,T].$ (8)
@@ 3.1 Reward fine-tuning in diffusion models
$\text{ s.t. }~\mathcal{S}(\lambda)={\operatorname*{arg\,min}_{p^{\prime}\in\mathcal{P}}}~\underbrace{-\mathbb{E}_{u\sim p^{\prime}}[r_{2}(u)]+\lambda\mathrm{KL}(p^{\prime}\|p_{\text{data}})}_{g(\lambda,p):=}$
@@ 3.2 Noise scheduling in diffusion models
$\text{ s.t. }~~\mathcal{S}(q)={\operatorname*{arg\,min}_{\theta^{\prime}\in\mathbb{R}^{d}}}~g(q,\theta):=\mathrm{L}_{\text{SM}}(\theta^{\prime},u(q))$ (10)
@@ 4.1 Strategy with pre-trained diffusion models
where $p_{\gamma}^{*}(\lambda)\in\operatorname*{arg\,min}_{p}\mathcal{L}_{\gamma}(\lambda,p)$ and $p^{*}(\lambda)\in\operatorname*{arg\,min}_{p}g(\lambda,p)$. Following the gradient-based bilevel method (Kwon et al., 2023; Shen et al., 2025b), a direct way to estimate $\nabla\mathcal{L}_{\gamma}^{*}(\lambda)$ is to use guided backward sampling in Algorithm 5 with (11a) and (11b) to obtain samples from $p^{*}(\lambda)$ and $p_{\gamma}^{*}(\lambda)$, and then compute the KL divergence from samples using kernel-based probability estimation. However, it has two drawbacks: 1) each $\lambda$ update requires guided backward sampling, which is computationally expensive; and 2) when the number of samples is less than the dimensionality of the data, the covariance matrix becomes rank-deficient, which makes kernel-based density estimation impossible.
@@
5:  estimate $\bar{\nabla}\mathcal{L}_{\gamma}^{*}(x_{k})$ with $y^{k}\approx y^{*}$ and $z^{k}\approx z^{*}$ in (4)
@@
4:  estimate $\bar{\nabla}\mathcal{L}_{\gamma_{k}}^{*}(\lambda_{k})$ by (33)
@@ 4.2 Strategy without pre-trained diffusion models
$\widetilde{\nabla}_{\theta}\mathrm{L}_{\text{SQ}}(\tilde{u}_{\theta,q}^{m})$ $=\frac{\xi}{2\nu}(\mathrm{L}_{\text{SQ}}(\tilde{u}_{\theta+\nu\xi,q}^{m})-\mathrm{L}_{\text{SQ}}(\tilde{u}_{\theta-\nu\xi,q}^{m}))$ (14)
@@
Parametrization for noise scheduler. To further reduce the memory cost and ensure the nondecreasing of $q(t)$, we parameterize the noise scheduler by the commonly used cosine and sigmoid function (Nichol & Dhariwal, 2021; Chen, 2023; Kingma et al., 2021) and optimize the parameters within these functions instead of directly optimizing $\{q(t)\}_{t=1}^{T}$. In both parametrization, we optimize just $4$ scalar parameters, significantly reducing the dimensionality of the optimization from $T$ to $4$. Specifically, we use
@@
where $q_{s},q_{e},q_{\epsilon},q_{\tau}$ represent the start, end, offset error, and power effect, respectively. Sigmoid parameterization is defined similarly in (34) and can be found in Appendix F.2. Since $q(t)$ should be nondecreasing, we assign q(t)=1l(t)/l(t1)q(t)=1-l(t)/l(t-1). After parameterization, ZO perturbation will be added on $q_{s},q_{e},q_{\epsilon},q_{\tau}$ instead of directly on $q(t)$.
@@ 5 Theoretical Guarantee
This theorem demonstrates that when $G_{\eta_{k},\gamma_{k}}(x_{k})\neq 0$, and setting penalty constant $\gamma_{k}\geq\frac{4\sqrt{3}B}{\|G_{\eta_{k},\gamma_{k}}(x_{k})\|}$, the bilevel algorithm will have strict descent over the hyper-function $F(x)$. Otherwise, since $\|\bar{\nabla}\mathcal{L}_{\gamma}^{*}(x)-\nabla F(x)\|={\cal O}(1/\gamma_{k})$, $G_{\eta_{k},\gamma_{k}}(x_{k})=0$ is approximately the stationary points of $F(x)$. The proof can be found in Appendix D.3.
@@
Grid search 125.77±1.3125.77\ $\scriptstyle\pm$ 1.3} 31.72±2.131.72\ $\scriptstyle\pm$ 2.1} $3.34$
Random search 123.70±4.3123.70\ $\scriptstyle\pm$ 4.3} 35.70±1.135.70\ $\scriptstyle\pm$ 1.1} $3.37$
Bayesian search 140.35±2.4140.35\ $\scriptstyle\pm$ 2.4} 33.29±2.933.29\ $\scriptstyle\pm$ 2.9} $12.7$
Weighted sum 116.52±3.9116.52\ $\scriptstyle\pm$ 3.9} 36.50±1.136.50\ $\scriptstyle\pm$ 1.1} $4.31$
@@
Table 1: Best FID and CLIP score given by different baselines and our method with penalty constant $\gamma=10^{3}$ for fine-tuning diffusion model application using synthetic lower-level reward (Yuan et al., 2024). Running time is measured in hours.
@@ 6 Numerical Experiments
In this section, we present the experimental results of the proposed bilevel-diffusion algorithms in two applications: reward fine-tuning and noise scheduling for diffusion models, and compare them with baseline hyperparameter optimization methods: grid search, random search, and Bayesian search (Snoek et al., 2012).
@@ 6.1 Reward fine-tuning in diffusion models
Furthermore, to showcase the robustness of our approach with respect to different reward functions at the lower level, we test the performance of each method for benchmarking reward function, HPSv2 (Wu et al., 2023), as the lower-level reward function. Comparisons of our method and different baselines are given in Table 2. Similar to the results given by the synthetic reward function (Yuan et al., 2024), bilevel approach also outperform other baselines in terms of image quality in comparable time complexity. We also provided the visualization of the generated images in Figure 5. While all HPO on fine-tuned models enhance the aesthetic, clarity and sharpness compared to the pre-trained image, the random, grid, Bayesian search, and weighted sum approaches messed up the legs and trunk, and fail to generate the right number of elephant’s legs. In comparison, our proposed bilevel approach not only generate colorful images, but also preserve correct elephant biological feature.
Baselines FID $\downarrow$ CLIP $\uparrow$ Time $\downarrow$
Grid search 142.76±2.4142.76\ $\scriptstyle\pm$ 2.4} 36.3±1.736.3\ $\scriptstyle\pm$ 1.7} $3.34$
Random search 139.21±3.1139.21\ $\scriptstyle\pm$ 3.1} 35.20±3.135.20\ $\scriptstyle\pm$ 3.1} $3.37$
Bayesian search 153.25±1.2153.25\ $\scriptstyle\pm$ 1.2} 34.98±2.934.98\ $\scriptstyle\pm$ 2.9} $12.7$
Weighted sum 140.56±1.3140.56\ $\scriptstyle\pm$ 1.3} 35.40±2.135.40\ $\scriptstyle\pm$ 2.1} $4.31$
Bilevel method $\textbf{137.23}\ {\scriptstyle\pm 1.7}$ $\textbf{37.10}\ {\scriptstyle\pm 3.2}$ $3.35$
Table 2: Best FID and CLIP score given by different baselines and our method with penalty constant $\gamma=10^{3}$ for fine-tuning diffusion model application using HPSv2 (Wu et al., 2023) as the lower-level reward. Running time is measured in hours.
Figure 5: Visualization of images generated by different methods using the prompt “elephant” and HPSv2 reward (Wu et al., 2023).
6.2 Noise scheduling in diffusion models
We evaluated our bilevel noise scheduling method, detailed in Algorithm 6, paired with DDIM backward sampling for the image generation on the MNIST dataset. We trained a U-Net with $178$ layers and $10^{6}+$ parameters following the github repository  [^1] https://github.com/bot66/MNISTDiffusion/tree/main. We considered both cosine and sigmoid parametrization and tuned $4$ parameters $q_{s},q_{e},q_{\tau},q_{\epsilon}$ jointly.
Table 3 presents the best FID, inception score (IS), and time complexity achieved by each method. The bilevel method achieves comparable performance with the hyperparameter optimization baselines in both FID and IS while being $6\times$ time faster. Generated images by each method are shown in Figure 7. While the Bayesian approach achieves relatively better FID and IS, it tends to focus on generating simpler numbers, such as $1$ and $7$. In contrast, the images produced by the bilevel method show excellent diversity across numbers 090-9 while maintaining image fidelity.
@@ 7 Related works
Fine-tuning diffusion models. Fine-tuning diffusion models aims to adapt pre-trained models to boost the reward on downstream tasks. Methods in this domain include directly backpropagating the reward (Clark et al., 2024), RL-based fine-tuning (Fan et al., 2024; Black et al., 2023), direct latent optimization (Tang et al., 2024; Wallace et al., 2023; Hoogeboom et al., 2023), guidance-based approach (Guo et al., 2024; Chung et al., 2022; Bansal et al., 2023) and optimal control (Uehara et al., 2024). Although entropy regularization is often incorporated into the reward to prevent over-optimization, no existing work has explored designing an efficient bilevel method to tune its strength.
Noise scheduling in diffusion models. Noise schedule is crucial to balance the computational efficiency with data fidelity during image generation. Early works, such as DDPM (Ho et al., 2020), employed simple linear schedules for noise variance, while Nichol & Dhariwal (2021) and Kingma et al. (2021) introduced cosine and sigmoid schedules to enhance performance. Recent studies (Lin et al., 2024; Chen, 2023) have highlighted limitations in traditional noise schedules and proposed new parameterization to improve the image quality. Notably, Sahoo et al. (2024) learned the noise scheduler by optimizing the log-likelihood, which yields a tighter lower bound (ELBO) and thus improves the generation quality of the diffusion model. However, none of the prior works considered using bilevel optimization to automatically learn the noise schedule for directly optimizing sample quality.
Bilevel hyperparameter optimization. Bilevel optimization has been explored as an efficient hyperparameter optimization framework, including hypernetwork search (Mackay et al., 2019; Liu et al., 2019), hyper-representation (Franceschi et al., 2018), regularization learning (Shaban et al., 2019) and data reweighting (Shaban et al., 2019; Franceschi et al., 2017). Recently, it has been explored in federated learning (Tarzanagh et al., 2022) and LLM fine-tuning (Shen et al., 2025a; Zakarias et al., 2024). None of the existing works have explored hyperparameter optimization in diffusion models, and the methods proposed so far are inapplicable due to the infinite-dimensional probability space and the high computational cost of sampling.